OpenStudy (anonymous):

A kindergarten school has square mats in two colors, red and blue. The ratio of the perimeter of the red mat to the perimeter of the blue mat is 2 : 1. The area of the blue mat is 36 square feet. What is the area of the red mat?

OpenStudy (anonymous):

What do you think I got 72 because I doubled it if the red perimeter is bigger

OpenStudy (diyadiya):

\[\text{ Side of Blue mat }= \sqrt{36}=6ft\]\[\text {Perimeter of Blue mat}= 4 \times 6 =24\]\[\frac{\text{perimeter of red mat}}{\text{perimeter of blue mat}}= \frac{2}{1}\]\[\frac{x}{24}= \frac{2}{1}\]\[x=2 \times 24 =48\] \[\text{Where x is perimeter of Red mat }\]\[\text{Side of Red mat }= \frac{48}{4}=12\]\[\text{Area of Red mat}= 12^2=144 sq.ft\]

OpenStudy (anonymous):

Sorry, this is the right answer.

OpenStudy (anonymous):

Since ou guys are viewing can you please look at my rebecca question